OpenStudy (cutepuppies):

Curtis drank 1/2 of his glass of water. Isa drank 2 1/3 times more water than Curtis. How many glasses of water did Isa drink? Express your answer in simplest form. A. 2 1/6 B. 1 1/6 C. 1 1/3 D. 6/7

OpenStudy (iluvsoccer):

Multiply both of them. First you have to convert the Whole Fraction thing. 2 x 3 + 1 is?

OpenStudy (jdoe0001):

make 2 1/3 to an "improper" fraction first with only 1 numerator and one denominator, then check how much is that times 1/2 :)

OpenStudy (study312):

it's called a improper fraction

OpenStudy (study312):

so what is 2 x 3?

OpenStudy (jdoe0001):

\(\bf 2\frac{1}{3}\iff \cfrac{\square ?}{\square ?}\times \cfrac{1}{2}\)

OpenStudy (iluvsoccer):

@cutepuppies, i'll help you. What is 2 x 3?

OpenStudy (study312):

nice pic @jdoe0001

OpenStudy (cutepuppies):

6

OpenStudy (iluvsoccer):

Now 6 + 1 is what?

OpenStudy (cutepuppies):

7

OpenStudy (iluvsoccer):

Good now our new numerator is 7 and we keep the same denominator. \[\frac{ 7 }{ 3 }\times \frac{ 1 }{ 2 }\]

OpenStudy (iluvsoccer):

now we just multiply across

OpenStudy (iluvsoccer):

7 x 1 is?

OpenStudy (cutepuppies):

7/6

OpenStudy (cutepuppies):

how do you simplify

OpenStudy (iluvsoccer):

Good now 7 divided by 6 is? Do long division. :)

OpenStudy (iluvsoccer):